OpenStudy (anonymous):

Calculate the first four terms of the following sequence, starting with n=1. Cn=5+(5/2)+(5/3)+(5/4).......(5/n).??? need some help

OpenStudy (amistre64):

hold that :)

OpenStudy (amistre64):

A{1} = 5/1 A{2} = 5 + 5/2 A{3} = 5 + 5/2 + 5/3 A{4} = 5 + 5/2 + 5/3 + 5/4 thats better

OpenStudy (amistre64):

5, 7.5, 9.1666.... , 10.416666......

OpenStudy (anonymous):

so for a problem like that you just take waht n = and whatever the last term is and go that far out with the terms:?

OpenStudy (anonymous):

like this one ... Calculate the first four terms of the following sequence, starting with n=1 b1=4 b2=2 bn=bn−1+4bn−2

OpenStudy (amistre64):

it depends on how they define the progression;

OpenStudy (amistre64):

what 'equation' do they give you to compute the terms by?

OpenStudy (amistre64):

b{1}=4; b{2}=2; b{n}=b{n−1} +4 b{n−2} ; thats easier to read

OpenStudy (anonymous):

its b sub (n-1) + 4b sub(n-2)

OpenStudy (amistre64):

b{1} = 1 b{2} = 2 b{3} = 2 +4(1) = 6 b{4} = 6 +4(2) = 14 b{5} = 14 +4(6) = 38

OpenStudy (anonymous):

not correct... the formula isnt b *(n-1) + 4 * b(n-2)...The (n-1) is a sub score to b same as (n-2)